OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3.

OpenStudy (anonymous):

can u do it

OpenStudy (anonymous):

parallet towat?

OpenStudy (anonymous):

to y=one thirdx -3

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. y = 3x - 14 y = one thirdx + 2 y = -3x + 22 y = -one thirdx + 8

OpenStudy (anonymous):

u have to use this equation (y-y1)=m(x-x1) m=slope the line's slope is equal to the given line's slope

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnaks

OpenStudy (anonymous):

ur welcome :)